OpenStudy (anonymous):

Evaluate this expression. P = 10, R = -4, T = 5 P 2 - R 2 =

OpenStudy (anonymous):

those 2's are suppose to be little like 10 to the second power

OpenStudy (anonymous):

Then you put ^2 for them

OpenStudy (anonymous):

plug in the numbers

OpenStudy (anonymous):

10(2)=20 -(-2)(2)=-4 20-4=16 :D